# Prediction-Enhanced Monte Carlo: A Machine Learning View on Control Variate

### v2

@@ 3.3. Prediction-Enhanced Monte Carlo: Prediction Model
    We then draw samples of $\boldsymbol{\theta}_{i}$ from $\boldsymbol{\Theta}$, using uniform sampling or any measure that ensures comprehensive coverage. For each sampled parameter set, we generate one pair  [^2] You could also generate multiple pair of $(\boldsymbol{Y}(\boldsymbol{\theta}_{i}),\boldsymbol{X}(\boldsymbol{\theta}_{i}))$ under the same sampled $\boldsymbol{\theta}_{i}$. of $\boldsymbol{Y}(\boldsymbol{\theta}_{i})$ and its corresponding $\boldsymbol{X}(\boldsymbol{\theta}_{i}):=\phi(\boldsymbol{Y}(\boldsymbol{\theta}_{i}))$. It’s important to note that $\boldsymbol{X}$ and $\boldsymbol{Y}$ are coupled in this generation process.
@@
    The machine learning model $g$ is then trained to minimize the mean squared error (MSE) loss and saved:
@@
Algorithm 1 Prediction Model Training
@@
This pre-training setup presents an ideal scenario for machine learning applications. It is important to also note that PEMC does not require $g$ to take specific forms. However, the convex and differentiable nature of the loss function (i.e., squared loss), the large supply of data (i.e., from Monte Carlo simulation), and the need for expressiveness of the model family make neural networks (NNs) an ideal choice. For training NNs, various optimization techniques and architectural enhancements—such as the Adam optimizer [80], batch normalization [75], and skip connections [64]—can be employed to improve performance. In fact, recent advancements in machine learning technologies have made the training process remarkably accessible and efficient. The availability of open-source, user-friendly frameworks like PyTorch [103], TensorFlow [1], and JAX [24]—coupled with GPU acceleration and highly optimized C++ backends—has dramatically simplified the process of training models on large datasets. These frameworks support recent neural network architectures, such as Convolutional Neural Networks (CNNs) [83], Residual Networks (ResNets) [64], and Transformers [128], as well as various Stochastic Gradient Descent (SGD) optimization algorithms [21]. In applications, we only utilized well-established and widely recognized neural network structures that are now considered common. These models can typically be implemented using just a few lines of code, as they have been incorporated into existing packages and extensively optimized.
Finally, regarding the choice loss function, we note that the goal of the prediction model $g$ is to produce an approximation of the underlying regression function:
$g(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}))$
$\approx$ $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}\bigg{[}f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}))\bigg{|}(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}))\bigg{]}$
As we will see, the quality of this approximation will impact the effectiveness of variance reduction, which we shall explore in detail in subsequent sections.
3.4. Prediction-Enhanced Monte Carlo: Evaluation
Having completed the training phase of PEMC, we now turn our attention to the evaluation procedure. This stage employs the pre-trained model to enhance the efficiency of Monte Carlo simulations, and we will demonstrate how control variates are cleverly constructed to achieve variance reduction. The evaluation process proceeds as follows:
@@ 3.4. Prediction-Enhanced Monte Carlo: Evaluation
    A specific parameter $\boldsymbol{\theta}:=(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}})\in\boldsymbol{\Theta}$ is given, e.g., calibrated from market data in real-time, and needs to be used for pricing.
@@
    Generate $n$ pairs of $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$, from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$. Here we have suppressed $\boldsymbol{Y}$’s dependency on $\boldsymbol{\theta}$, as in the evaluation stage only one fixed $\boldsymbol{\theta}$ is considered.
@@
    Based on the marginal distribution of $\boldsymbol{X}$, independently generate $N$ additional samples of $(\boldsymbol{\tilde{X}}_{j})_{j\in[N]}$ from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$, directly according to its marginal distribution. Importantly, these samples are independent of the previous data $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$.
@@
Algorithm 2 Evaluation
3.5. PEMC: An Example on Asian Options
We now return to our example on arithmetic Asian option under the Heston model:
$P_{A}(\{S_{t}\}_{t})=\max\left(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0\right)$
where $\boldsymbol{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$. The Heston model is described in (1) which requires $\boldsymbol{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$. For this example, let’s say we use a neural network (NN) as the prediction model. Then, the PEMC framework could proceed as follows:
@@ 3.5. PEMC: An Example on Asian Options
    At evaluation, a specific $\boldsymbol{\theta}=({r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu})\in\Theta$ is given, we generate $n$ paired samples $(\textit{label}_{i},\textit{feature}_{i})_{i\in[n]}$ as in (4). We also generate $N$ i.i.d. $\sqrt{T}\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$ samples and store them as $(\tilde{\textit{feature}})_{j\in[N]}$.
@@
3.6. Implementation Details in PEMC
Having walked through the Asian option example under the PEMC framework, we discuss several important implementation details. These considerations are critical for successfully applying PEMC in practice.
3.6.1. Parameter Space $\Theta$
3.6.2. The Design of $\Large{X}$: Feature Engineering
In the example, we designed the feature $\boldsymbol{X}$ as the sum of two correlated Brownian increments. As discussed in Section 3.3, the choice was motivated by two key properties:
  • Predictability: The sum of Brownian motion increments explains away a portion of the variance of the SDE.
Given these considerations, it is tempting to employ Gaussian sums as a generic recipe for the majority of option pricing problems handled by PEMC, particularly since these problems commonly involve stochastic differential equations driven by Brownian motion. Indeed, this is precisely the approach we adopt in our applications, demonstrating its robustness and effectiveness. In cases involving long simulation paths, one might consider segmenting the sum into smaller increments, though we observed that a single increment usually suffices to achieve significant variance reduction. However, the choice of Brownian increments, while effective, is certainly not the only one. As with most ML tasks, feature engineering that leverages domain expertise can yield better results. We illustrate with one such example, the pricing of a floating strike lookback call option, under the Heston model(1). The option payoff is given by:
$f(\{S_{t}\}_{t})=\max(S_{T}-\min_{0\leq t\leq T}S_{t},0)$
The rationale behind the construction of $\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W_{t}^{S}$ term is grounded in the domain knowledge of the Heston model. Specifically, one who is familiar with the role of $\eta$ as the mean-reverting level of variance process $\nu_{t}$ may approximate $\nu_{t}\approx\eta$ (i.e., assume fixed volatility) and simplify the asset price process to a geometric Brownian motion:
This approximation leads to the expression: $f(\{S_{t}\}_{t})\approx S_{0}e^{(r-\nu/2)T+\sqrt{\nu}[\boldsymbol{X}]_{1}}-S_{0}e^{[\boldsymbol{X}]_{2}}$, which could serve as a potentially effective approximation of $f(\{S_{t}\}_{t})$. This corresponds directly to Property 1 in the construction of $\boldsymbol{X}$ in 3.3. For Property 2, the marginal distribution of $\boldsymbol{X}$ permits efficient and parallelizable simulation. In particular, applying Girsanov’s Theorem enables one to derive a closed-form expression for the joint density of $([\boldsymbol{X}]_{1},[\boldsymbol{X}]_{2})$ [78]. This result makes it straightforward to sample these components directly.
3.6.3. Eliminating Data Storage Overheads
In typical ML workflows, training datasets—often costly, scarce, and carefully curated—are stored and reused extensively. By contrast, within the PEMC framework, the nature of data generation during the prediction model training in Algorithm 1, actually does not require extensive data storage. This difference stems from two main considerations. First, the volume of training data required to achieve a well-performing prediction model can be large. Storing all of it would be both expensive and unnecessary. In PEMC, the training data can be produced directly via Monte Carlo simulation, ensuring an effectively unlimited supply. Second, this flexibility allows for a more efficient workflow. Data can be generated “on the fly” and processed in streaming fashion. For example, to train on a large number of $N_{\text{train}}$ samples, one could iteratively produce small batches, train the model on these batches, and discard them after training. In this manner, we have successfully trained models in the applications using data on the order of $10^{7}\textendash 10^{8}$.
@@ 4. Analysis and Performance
In this section, we provide a detailed and comprehensive analysis of the PEMC estimator. All properties of PEMC discussed here pertain to the evaluation stage, where $\boldsymbol{\theta}$ is arbitrary but fixed. Henceforth, we omit the explicit dependence on $\boldsymbol{\theta}$ in the notation (e.g., writing $\boldsymbol{Y}$ in place of $\boldsymbol{Y}(\boldsymbol{\theta})$). We begin by establishing the unbiasedness of our estimator.
@@ 5.3.3. Ambulance Diversion Policies Evaluation
The simulation workflow spans nested simulation system. We focus here on a representative subproblem—the two-hospital ambulance diversion— to illustrate PEMC’s advantage. The discrete-event simulation models two emergency departments, ED-1 and ED-2, each with physicians, patients, and arrival streams whose rates $\lambda_{h,d}$ vary by hour $h$ and day $d$. Ingredients include nonhomogeneous Poisson process with hourly arrival/service rates $\lambda_{h,d}$, e.g., $\text{Arrivals}(t)\sim\text{NHPP}\bigl{(}\lambda(t)\bigr{)}$; Triage levels and service times where each patient is assigned a triage level $\ell\in\{1,2,3,4,5\}$ via a (possibly crisis-adjusted) multinomial distribution; Priority queueing and threshold diversion where physicians are modeled as resources and patients forms priority queue ordered by triage. A threshold $\tau$ controls ambulance diversion: if the queue $Q$ in one hospital exceeds $\tau$, new ambulance arrivals divert to the other hospital, incurring additional travel time $\Delta$ and so on. The evaluation needed is how weekly mortality counts depends on the diversion threshold $\tau$. By varying $\tau$ in the simulation, we obtain mortality counts (or rates) as a function of policy choices, enabling downstream decisions. For PEMC predictor we employ a random forest estimator—well-suited for context [35, 105]. We gather a training size $10^{5}$, and sample $\boldsymbol{\theta}$ from a $\boldsymbol{\Theta}$ calibrated from EMS data. The $\boldsymbol{X}$ are readily simulated components for the nested MC—drawing on Weibull, Exponential, Gamma, and Poisson distributions with $\dim(\boldsymbol{X})=12$and is also parallelizable. The results are summarized Figure 5 and Table7. Further details are left in Appendix E.
@@ 6. Conclusion
Appendix
Appendix A Learning Theory Proofs
@@
Appendix B Approximation Error $\epsilon_{a}^{\mathcal{G}}$
@@
Appendix C Statistical Error $\epsilon_{e}^{N_{\text{train}}}$
@@ Appendix D Proof of Lemma 6
D.1. Variance Reduction of PEMC: An Example on Asian Options
@@ D.1.2. Evaluation
In particular, when a control variate method is available, such as in our current example, PEMC can be built upon it to potentially achieve even greater variance reduction. To see how, in Algorithm 1, we change the label from $P_{A}(\{S_{t}\}_{t})$ (MC baseline) to $P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})+P_{G}^{\text{exact}}(\boldsymbol{\theta})$ (Geometric CV baseline), and let the model $g$ learn
@@
Appendix E Ambulance Diversion
Details can be further found in [41, 45]. Some key ingredients are 1. Patient Arrivals: We simulate patient inflow to each hospital (A or B) according to a nonhomogeneous Poisson process with hourly rates $\lambda_{h,d}$. Specifically:
@@ Appendix E Ambulance Diversion
where $\lambda(t)$ is determined by a piecewise schedule over each hour $h$ of day $d$. Many of these arrivals (e.g., 25%) come by ambulance, with the remainder as walk-ins. 2. Triage Levels and Service Times: Each arriving patient is assigned a triage level $\ell\in\{1,2,3,4,5\}$ via a (possibly crisis-adjusted) multinomial distribution:
@@
4. Mortality Function Each patient $i$ has a time-varying risk of death. The code uses two complementary representations: 1. Direct Mortality Probability:
@@
$\min_{\tau}\;\mathbb{E}\Bigl{[}\sum_{\text{patients }i}\mathbf{1}\bigl{\{}\text{wait}_{i}>\text{death\_time}_{i}\bigr{\}}\Bigr{]},$

### v3

@@ 3.3. Building Prediction Model
    We then draw samples of $\boldsymbol{\theta}_{i}$ from $\boldsymbol{\Theta}$, using uniform sampling or any measure that ensures comprehensive coverage. For each sampled parameter set, we generate one pair  [^3] You could also generate multiple pair of $(\boldsymbol{Y}(\boldsymbol{\theta}_{i}),\boldsymbol{X}(\boldsymbol{\theta}_{i}))$ under the same sampled $\boldsymbol{\theta}_{i}$. of $\boldsymbol{Y}(\boldsymbol{\theta}_{i})$ and its corresponding $\boldsymbol{X}(\boldsymbol{\theta}_{i}):=\phi(\boldsymbol{Y}(\boldsymbol{\theta}_{i}))$. It is important to note that $\boldsymbol{X}$ and $\boldsymbol{Y}$ are coupled in this generation process.
@@
    The ML model $g$ is then trained to minimize the MSE loss and saved:
@@
Algorithm 1 Prediction Model Training in PEMC
@@
This pre-training setup presents an ideal scenario for ML applications. It is important to also note that PEMC does not require $g$ to take specific forms. However, the convex and differentiable nature of the loss function (i.e., squared loss), the large supply of data (i.e., from Monte Carlo simulation), and the need for expressiveness of the model family make neural networks (NNs) an ideal choice. For training NNs, various optimization techniques and architectural enhancements—such as the Adam optimizer [81], batch normalization [76], and skip connections [65]—can be employed to improve performance. In fact, recent advancements in ML technologies have made the training process remarkably accessible and efficient. The availability of open-source, user-friendly frameworks like PyTorch [105], TensorFlow [1], and JAX [24]—coupled with GPU acceleration and highly optimized C++ backends—has dramatically simplified the process of training models on large datasets. These frameworks support recent neural network architectures, such as Convolutional Neural Networks (CNNs) [84], Residual Networks (ResNets) [65], and Transformers [131], as well as various Stochastic Gradient Descent (SGD) optimization algorithms [21]. In applications, we only utilized well-established and widely recognized NN structures that are now considered common. These models can typically be implemented using just a few lines of code, as they have been incorporated into existing packages and extensively optimized.
3.4. Evaluation
Once we finish the pre-training phase, the prediction model is stored and used on-the-fly in the evaluation process as follows:
@@ 3.4. Evaluation
    A specific parameter $\boldsymbol{\theta}:=(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}})\in\boldsymbol{\Theta}$ is given, e.g., calibrated from market data in real time, and needs to be used for pricing.
@@
    Generate $n$ pairs of $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$, from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$. Here we have suppressed $\boldsymbol{Y}$’s dependency on $\boldsymbol{\theta}$ for convenience.
@@
    Based on the marginal distribution of $\boldsymbol{X}$, independently generate $N$ additional samples of $(\boldsymbol{\tilde{X}}_{j})_{j\in[N]}$ from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$, directly according to its marginal distribution. These samples are independent of the previous data $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$.
@@
Algorithm 2 Evaluation Procedure in PEMC
3.5. Overall Methodology
We now summarize our overall PEMC methodology applied to the Asian option under the Heston model, i.e., (7), with $\boldsymbol{\theta_{\text{payoff}}}$ specifying the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$. The Heston model is described in (1) which requires $\boldsymbol{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$. Suppose we use an NN as the prediction model. Then, the PEMC framework could proceed as follows:
@@ 3.5. Overall Methodology
    At evaluation, given a specific $\boldsymbol{\theta}=({r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu})\in\Theta$, we generate $n$ paired samples $(\textit{label}_{i},\textit{feature}_{i})_{i\in[n]}$ as in (4). We also generate $N$ i.i.d. $\sqrt{T}\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$ samples and store them as $(\tilde{\textit{feature}})_{j\in[N]}$.
@@
In Appendix B, we present numerical results on the performance comparisons of PEMC with standard MC and CV for Asian options under GBM. In this case, an excellent CV using the Black-Scholes formula is available under the standard CV framework for us to benchmark against. At the same time, in Section 5, we present results on more realistic and technically complex applications, including several production-grade exotic options that are vastly more complex than Asian options.
@@ 4. Analysis and Performance
In this section, we provide a detailed and comprehensive analysis of the PEMC estimator. All omitted proofs are delegated to Appendix A. All properties of PEMC discussed here pertain to the evaluation stage, where $\boldsymbol{\theta}$ is arbitrary but fixed. Henceforth, we omit the explicit dependence on $\boldsymbol{\theta}$ in the notation (e.g., writing $\boldsymbol{Y}$ in place of $\boldsymbol{Y}(\boldsymbol{\theta})$).
@@ 5.3.3. Ambulance Diversion Policies Evaluation
The simulation workflow spans nested simulation system. We focus here on a representative subproblem—the two-hospital ambulance diversion— to illustrate PEMC’s advantage. The discrete-event simulation models two emergency departments, ED-1 and ED-2, each with physicians, patients, and arrival streams whose rates $\lambda_{h,d}$ vary by hour $h$ and day $d$. Ingredients include nonhomogeneous Poisson process with hourly arrival/service rates $\lambda_{h,d}$, e.g., $\text{Arrivals}(t)\sim\text{NHPP}\bigl{(}\lambda(t)\bigr{)}$; Triage levels and service times where each patient is assigned a triage level $\ell\in\{1,2,3,4,5\}$ via a (possibly crisis-adjusted) multinomial distribution; Priority queueing and threshold diversion where physicians are modeled as resources and patients forms priority queue ordered by triage. A threshold $\tau$ controls ambulance diversion: if the queue $Q$ in one hospital exceeds $\tau$, new ambulance arrivals divert to the other hospital, incurring additional travel time $\Delta$ and so on. The evaluation needed is how weekly mortality counts depends on the diversion threshold $\tau$. By varying $\tau$ in the simulation, we obtain mortality counts (or rates) as a function of policy choices, enabling downstream decisions. For PEMC predictor we employ a random forest estimator—well-suited for context [35, 108]. We gather a training size $10^{5}$, and sample $\boldsymbol{\theta}$ from a $\boldsymbol{\Theta}$ calibrated from EMS data. The $\boldsymbol{X}$ are readily simulated components for the nested MC—drawing on Weibull, Exponential, Gamma, and Poisson distributions with $\dim(\boldsymbol{X})=12$and is also parallelizable. The results are summarized Figure 5 and Table7. Further details are left in Appendix C.
@@ 6. Conclusion
Appendix A Proofs
A.1. Proof of Lemma 3
Proof.
s.t. $nc_{f-g}+Nc_{g}\leq B,$
A.2. Proof of Lemma 4
Proof.
A.3. Proof of Lemma 5
Proof.
and $f$ is square-integrable, it follows that $\mathbb{E}[(f-g)h]=0$ for all square-integrable $h$. Plugging in $h=g$, we obtain $\mathrm{Cov}(f-g,g)=0$. Consequently we have $\rho:=\mathrm{corr}(f,g)=\frac{\sigma_{g}}{\sigma_{f}}$ and $\sigma_{f}^{2}=\sigma_{f-g}^{2}+\sigma^{2}_{g}$, which further implies $\sigma_{g}=\rho^{2}\sigma_{f}^{2}$ and $\sigma_{f-g}^{2}=(1-\rho^{2})\sigma^{2}_{f}$. The rest follows from $c_{f-g}=c_{f}$. ∎
A.4. Proof of Lemma 6
@@
A.4.1. Approximation Error $\epsilon_{a}^{\mathcal{G}}$
@@
A.4.2. Statistical Error $\epsilon_{e}^{N_{\text{train}}}$
@@ A.4.3. Proof of Lemma 6
Appendix B Numerical Results on Asian Option Pricing
@@ B.2. Evaluation
B.3. Using PEMC as a “Boost" to Known CV
Lastly, we demonstrate how PEMC can be used to improve existing CV. This supports the creative use of PEMC for further efficiency enhancement even for problems that already possess some variance reduction approaches. In particular, when a CV with known mean is available, such as in this current example, PEMC can be built upon it to potentially achieve even greater variance reduction. To see how, in Algorithm 1, we change the label from $P_{A}(\{S_{t}\}_{t})$ (MC baseline) to $P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})+P_{G}^{\text{exact}}(\boldsymbol{\theta})$ (Geometric CV baseline), and let the model $g$ learn
@@ B.3. Using PEMC as a “Boost" to Known CV
Appendix C Additional Details on the Ambulance Diversion Example
We provide more details on the ambulance diversion policy evaluation example in Section LABEL:sec:EMS. Some key ingredients are as follows: 1) Patient Arrivals: We simulate patient inflow to each hospital (A or B) according to a nonhomogeneous Poisson process with hourly rates $\lambda_{h,d}$. Specifically:
@@ Appendix C Additional Details on the Ambulance Diversion Example
where $\lambda(t)$ is determined by a piecewise schedule over each hour $h$ of day $d$. Many of these arrivals (e.g., 25%) come by ambulance, with the remainder as walk-ins. 2) Triage Levels and Service Times: Each arriving patient is assigned a triage level $\ell\in\{1,2,3,4,5\}$ via a (possibly crisis-adjusted) multinomial distribution:
@@
4) Mortality Function Each patient $i$ has a time-varying risk of death. The code uses two complementary representations: 1. Direct Mortality Probability:
@@
$\min\;\mathbb{E}\Bigl{[}\sum_{\text{patients }i}\mathbf{1}\bigl{\{}\text{wait}_{i}>\text{death\_time}_{i}\bigr{\}}\Bigr{]},$
@@
Appendix D Further Implementation Details of PEMC
We discuss several further implementation details that are important for successfully applying PEMC in practice.
D.1. Parameter Space $\Theta$
D.2. Eliminating Data Storage Overheads
In typical ML workflows, training datasets—often costly, scarce, and carefully curated—are stored and reused extensively. By contrast, within the PEMC framework, the nature of data generation during the prediction model training in Algorithm 1, actually does not require extensive data storage. This difference stems from two main considerations. First, the volume of training data required to achieve a well-performing prediction model can be large. Storing all of it would be both expensive and unnecessary. In PEMC, the training data can be produced directly via MC simulation, ensuring an effectively unlimited supply. Second, this flexibility allows for a more efficient workflow. Data can be generated “on the fly” and processed in streaming fashion. For example, to train on a large number of $N_{\text{train}}$ samples, one could iteratively produce small batches, train the model on these batches, and discard them after training. In this manner, we have successfully trained models in the applications using data on the order of $10^{7}\textendash 10^{8}$.
D.3. Evaluation Metrics
While training the NN estimator $g$ using MSE loss is common, it is not always clear how to interpret the resulting MSE score. Unlike some well-established benchmarks (e.g., classification accuracy), there is no canonical threshold or known “good” MSE value for a given problem. This ambiguity makes it challenging to determine when the network is sufficiently trained. To address this, we can exploit the fact that $g$ is meant to represent the conditional expectation $g=\mathbb{E}[f\mid\text{input}]$. If the network approximates this expectation well, then the sample average of $g(\boldsymbol{X})$ should be close to the sample average of $f(\boldsymbol{Y})$ over a given dataset. One practical diagnostic is to compute the Mean Absolute Relative Error (MARE) between these two averages. If $\mathbb{E}[g(\boldsymbol{X})]\approx\mathbb{E}[f(\boldsymbol{Y})]$, it provides a tangible indication that $g$ is capturing the underlying expectation. Our empirical experience suggests this criterion is very effective in practice (a 5-1% MARE typically indicates exceptional PEMC), complementing common techniques like early stopping in ML workflows. More importantly, as the we have shown in the theory, if $g$ closely approximates the conditional expectation, the variance reduction in PEMC is guaranteed—even if marginally—relative to standard MC.
Appendix E Additional Details on Hyper-Parameters for Variance Swaps
Tables 12 and 13 detail the hyper-parameter setups for the NNs presented in Section LABEL:sec:var_swap.
Table 12. Neural Network Architecture Parameters
CNN Branch Feed-forward Branch
kernel size: 3, stride: 1, padding: 1
max pooling (kernel: 2, stride: 2, padding: 0)
hidden dim: 512
output dim: 128
Feed-forward Synthesizer Parameters
Hidden dim: 128
Output dim: 1
Table 13. Hyper-parameter setup for the neural network
Appendix F Review of HJM
The HJM model directly describes the evolution of the entire forward rate curve, offering greater flexibility than traditional short-rate models. For illustration purposes, we focus on a one-factor specification with exponential volatility structure [61], though the framework readily extends to multi-factor cases.
Consider the problem of pricing a swaption. This is a contract granting its holder the right, but not the obligation, to enter into an interest rate swap at a future date. In a standard interest rate swap, one party agrees to pay a fixed rate while receiving a floating rate, and the other party does the opposite. Consider a swap with $n_{p}$ fixed payment periods, each of length $\Delta t^{\prime}$, starting at time $t^{\prime}_{0}$ and ending at time $t^{\prime}_{n_{p}}=t^{\prime}_{0}+\sum_{l=1}^{n_{p}}\Delta t^{\prime}$. The value of this swap at time $t^{\prime}_{0}$ is:
where $C$ is the notional amount (contract size), $R$ is the fixed rate, and $B(t^{\prime}_{0},t^{\prime}_{l})$ is the discount factor from $t^{\prime}_{0}$ to $t^{\prime}_{l}$. A swaption provides the holder with the option to enter into this swap at $t^{\prime}_{0}$. The payoff of the swaption is simply
$\max(0,V_{t^{\prime}_{0}}),$
[…]
$df(t,T)=\mu(t,T)\,dt+\sigma(t,T)^{\top}dW(t),$
where $\mu(t,T)$ is the drift, $\sigma(t,T)$ is the volatility function of the forward rate, and $W(t)$ is a Brownian motion. In contrast to short-rate models (e.g., [130] or [39]), which only model the dynamics of the short-term interest rate, the HJM model directly models the dynamics of the entire term structure of interest rates [66]. The HJM model is widely used in practice because of its flexibility in modeling interest rate derivatives like swaptions and its ability to incorporate complex volatility structures [29, 5]. However, the model’s generality also leads to the need for sophisticated numerical methods for simulation [61]. A key property of the HJM model is the no-arbitrage condition [61], which specifies the drift completely by the volatility:
[…]